OpenStudy (anonymous):

Quick One: What's special about the numbers 8,589,934.592 and 116,415,321,826,934,814,453,125 ?

OpenStudy (anonymous):

Is that a decimal point in the first number?

OpenStudy (anonymous):

Oops, screw up, s/b comma sorry....

OpenStudy (anonymous):

IS IT A SPECIAL NUMBER?

OpenStudy (anonymous):

What's a special number?

OpenStudy (anonymous):

WELL I think 89 is the special number and I think u know why.

OpenStudy (anonymous):

It's on Wikipedia http://en.wikipedia.org/wiki/89_(number)

OpenStudy (anonymous):

what about 90

OpenStudy (anonymous):

I don't know about 90...

OpenStudy (anonymous):

It will be soon special

OpenStudy (anonymous):

Wow, I thought this would be answered in no time...

OpenStudy (anonymous):

@sauravshakya Come on, I will help u to 90:-)

OpenStudy (anonymous):

At first, I thought they might be Fibonacci numbers, but I checked and they aren't. Then I thought they might be from the cousins of Fibonacci numbers athat start with 1, 3, etc, but I can't remember the name of that series.

OpenStudy (anonymous):

BOTH ARE RATIONAL NUMBERS.......lol

OpenStudy (anonymous):

Not special enough...

OpenStudy (anonymous):

And both don't fit in my calculator.

OpenStudy (anonymous):

Um, maybe that counts....:-)

OpenStudy (anonymous):

The second one looks like our national debt.

OpenStudy (anonymous):

OK, I will give you a clue, the second one is 5^33

OpenStudy (anonymous):

There are no 0's in both numbers and when you multiply them you get 1000000000000000000000000000000000

OpenStudy (anonymous):

2^33 for the other

OpenStudy (anonymous):

10^(33) largest known power of 10 that can be represented by the product of 2 numbers that contain no zeros

OpenStudy (anonymous):

Something about no shared numbers between composites and its primes I think

OpenStudy (anonymous):

Isn't this a rather silly question? First, you need special software to work with numbers this large.

OpenStudy (anonymous):

No thats wrong cuz the first is not prime

OpenStudy (anonymous):

I just multiplied and guessed. Good question @estudier

OpenStudy (anonymous):

"Isn't this a rather silly question? First, you need special software to work with numbers this large." That's what experimental number theorists do for a living :)

OpenStudy (anonymous):

HOW TO MULTIPLY SUCH LARGE NUMBERS.

OpenStudy (anonymous):

Is that what you do?

OpenStudy (anonymous):

Yes, but they have special software. Here, students have calculators.

OpenStudy (anonymous):

Certainly not, I just read a lot...

OpenStudy (anonymous):

I would like to learn it.

OpenStudy (anonymous):

"Here, students have calculators." and Wolfram

OpenStudy (anonymous):

ahh wolfram. My lifeline to advanced math

OpenStudy (anonymous):

I'm going to go do something relevant and meaningful.

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

2^33 * 5^33 = (2*5)^33=10^33

OpenStudy (anonymous):

BUT HOW? 116,415,321,826,934,814,453,125=5^33

OpenStudy (anonymous):

It is conjectured that there is no larger power of 10 in similar way (or at least, we will never find it if there is)